# Minds versus Machines: Rethinking Entailment Verification with Language Models

### v1

@@ 3 Training LLMs for Entailment Verification
In Section 2.4, we observed that open-sourced LLMs are lacking in performance compared to humans and close-sourced models such as GPT-4. Thus, in this section, we propose training strategies to finetune an open-sourced model that can perform competitively w.r.t. GPT-4. To this end, we finetune Flan-T5-xxl model using the training data from some of the above-mentioned datasets. Specifically, we choose the training portions of ANLI, RACE, and ECQA for finetuning the model. Please refer to Appendix F.1 for more details on our training dataset selection criteria. We describe two approaches to finetuning and then discuss our findings.
@@
3.1 Finetuning Formulations
This section describes the two fine-tuning formulations explored in this work.
Classification
@@ 3.1 Finetuning Formulations
In this approach, for a given premise and hypothesis pair $(p,h)$, we define a weaker hypothesis $h^{\prime}$ as a statement such that the premise $p$ supports $h$ more strongly than $h^{\prime}$. Then, for a given triple $(p,h,h^{\prime})$, we formulate the ranking task as predicting the hypothesis that is more supported by the premise. Given the triplet $(p,h,h^{\prime})$, we define the margin ranking loss as follows:
@@
where $s(p,h)$ is the entailment score as defined in Equation 1. The key advantage of this formulation over the classification is that ranking, by design, learns a softer decision boundary between the two labels. This can lead to better generalization, especially for contextual QA datasets. Sometimes, the wrong choice can be relatively less favorable w.r.t. to the best option in QA instead of being absolutely incorrect. Training to hard-classify the hypothesis for such options can be avoided by ranking them with the best hypothesis (corresponding to the right choice), thus learning a softer classification boundary. We refer to the finetuned model using the ranking objective as “Flan-T5-xxl + Rank”. Please refer to Appendix F.2 for more details on the training data collection process for ranking.
@@ 3.2 Findings
Table LABEL:tab:finetune_results shows the evaluation results on the complete evaluation set (i.e., we use all the data points instead of 100 samples per dataset, which was used in Table LABEL:tab:human_results). For our models, we separately average the results for the datasets already seen in training (namely, ANLI, RACE, and ECQA) and unseen during training into two columns, seen and unseen, respectively. First, we observe that all our finetuned models are consistently better across nine out of ten datasets than the baseline Flan-T5-xxl. Finetuning improves 0.07 macro-F1 on average. This shows that finetuning is overall beneficial in training the model on the task of entailment verification.
@@
4 Application: Filtering CoT Rationales
Recently, Wang et al. (2023) proposed self-consistency (SC), a decoding technique to improve over chain-of-though (CoT) reasoning Wei et al. (2022) in LLMs, whereby multiple CoT rationales are sampled for a given input instance and a majority voting overall predicted labels is considered as the final prediction. But generative LLMs can potentially output rationales that are inconsistent Ye and Durrett (2022), i.e., the rationale does not support the corresponding model prediction. Such inconsistency can, in turn, degrade the overall self-consistency results. This motivates an important application of entailment verification systems in this paradigm. Please refer to Appendix G.1 for examples of consistent and inconsistent CoTs.
Approach
As shown in Figure LABEL:fig:cot_application, we can use a verifier as an intermediate step before computing the majority vote, whereby the inconsistent rationales are filtered out. We achieve this by first scoring all the rationales (premises) and their corresponding predictions (hypothesis [^2] For MCQ datasets, typically, the prediction is one of the choices. We use Chen et al. (2021) as defined in Section 2.1 to convert the question-answer pair into a statement.) using Equation 1. Then, we use the top-$k$ rationales for majority voting. We set $k=10$ for all our experiments.
@@ 4 Application: Filtering CoT Rationales
In figure LABEL:fig:cot_comparison, we compare the vanilla SC with the filtering+SC approach described above. Following Wang et al. (2023), we compute the average performance of these methods across three MCQ datasets for four different base CoT models (UL2 Tay et al. (2023), Codex-001 Brown et al. (2020), LaMDA-137B Thoppilan et al. (2022), and ChatGPT OpenAI (2022)). Please refer to Appendix G for details on the datasets and more comparisons with Flan-T5-xxl. We observe that filtering leads to a consistent performance gain over SC across all CoT base models. This demonstrates the advantage of the filtering approach. Next, we find that the improvements are more prominent for weaker base models such as UL2 than the stronger ones (ChatGPT). For instance, filtering UL2 generated-rationales can even achieve comparable performance with vanilla SC over LaMDA-137B. In comparison, the gains for filtering ChatGPT CoTs are $\sim$ 1.5 %. This shows that weaker models are prone to generating inconsistent CoTs and thus benefit more from this approach. But at the same time, even stronger models such as ChatGPT can still benefit from consistency checks. Please refer to Appendix G.1 for examples of filtered CoT rationales.
@@ 5 Related Works
With the advent of strong, general-purpose LLMs Brown et al. (2020); Chung et al. (2022); OpenAI (2023), many prompting strategies have been proposed to generate a natural language reasoning along with the model’s prediction Wei et al. (2022); Zhou et al. (2023); Yao et al. (2023); Huang and Chang (2023). Recently, Ye and Durrett (2022) have found that such generations can sometimes be unreliable due to non-factual and inconsistent reasoning, while Huang et al. (2023) have argued that LLMs struggle to self-correct such issues without external feedback.
Prior works have addressed this limitation by oversampling reasoning chains and marginalizing Wang et al. (2023), using the LLMs itself to recheck their reasoning Madaan et al. (2023); Miao et al. (2023), leveraging external knowledge source to verify factuality Zhao et al. (2023), using deterministic solvers to improve faithfulness Lyu et al. (2023), decomposing the reasoning steps into smaller steps Ling et al. (2023), etc. While the progress is impressive, some of these are either specialized approaches for math-specific datasets or heavily rely on close-sourced LLMs (GPT-3.5, GPT-4, etc.) for verification. In contrast, here we focus on natural language datasets and develop a strong open-sourced LLM that can be easily deployed to check entailment errors in LLM reasoning chains.
@@ 6 Conclusion
We studied the entailment verification problem in the context of LLMs. Specifically, we sourced datasets across three different categories (NLI, contextual QA, and rationales) and analyzed both human and model performance across these datasets. We found some misalignments between the two, whereby models are better than humans in complex reasoning and humans are relatively more consistent on simpler reasoning tasks. We also explored different finetuning objectives to train LLMs for verification. Our fine-tuned models outperform GPT-3.5 and are at par with GPT-4. Finally, we demonstrated a practical application of entailment verification in filtering out inconsistent model-generated rationales while using self-consistency decoding. Overall, we believe our work provides a meaningful comparison of inference capabilities between humans and LLMs, a comprehensive evaluation benchmark, an open-sourced entailment verification model, and interesting applications of the task.
@@
Even though our work demonstrates exciting results on entailment verification tasks by finetuning LLMs, several limitations can be potentially improved. We only tried encoder-decoder-based models for finetuning. However, other models with different architectures (like decoder-only) can also be considered. Our strategy to convert a QA pair into a statement using the QA-to-statement converter model can have errors that can cascade both in the evaluation dataset and our fine-tuned models. For computing the entailment score in Equation 1, we only considered the probability of “Yes” and “No” tokens in the entire vocabulary. Other alternative expressions like “YES”/“NO”, “True”/“False”, etc., can also be considered to make the score more robust. Finally, our training objective outputs entailment scores instead of directly generating answers. Answer generation as a training objective can be more robust since it is a stricter objective than our scoring technique.
@@ Acknowledgments
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200006, the DARPA MCS program under Contract No. N660011924033, the Defense Advanced Research Projects Agency with award W911NF-19-20271, NSF IIS 2048211, and gift awards from Google and Amazon. The views and conclusions contained herein are those of the authors. They should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. We would like to thank all the USC INK research lab collaborators for their constructive feedback on the work.
@@ Appendix A Evaluation Datasets
NLI is an obvious choice of data source as it is a more general case of the entailment verification problem. While converting an NLI dataset for our task, we merge the neutral and contradict labels to the not support label. We use the following NLI datasets in our benchmark:
  • WaNLI Liu et al. (2022): This is a new NLI dataset built using worker and AI collaboration. This challenging dataset improves over the existing NLI dataset MultiNLI Williams et al. (2018). And the test split is used when doing evaluation.
  • FEVER Nie et al. (2019): This is a modification of the original FEVER dataset Thorne et al. (2018) in which the claim is paired with textual evidence from Wikipedia to convert it into an NLI format dataset. This pairing uses existing state-of-the-art evidence extraction systems to find relevant evidence for each claim. Premises in this dataset typically contain multiple sentences, which is one of our focus areas. As the test split is not available, we report results on dev split for evaluation.
@@
Next, we consider QA datasets where the task is to answer a question based on a given context and some options. We use an off-the-shelf QA-to-statement converter model Chen et al. (2021) to generate a hypothesis statement for each question option pair. Then, the hypothesis corresponding to the correct choice is marked as “support”, while the rest are marked as “not support” to create the entailment verification dataset. Overall, we include the following datasets from this category:
  • Cosmos QA (CosQA) Huang et al. (2019): This dataset contains multiple-choice questions (MCQs) that require an understanding of commonsense-based reading comprehension to answer a question. The key challenge in this dataset is understanding people’s everyday narratives described in the context that can have some missing commonsense knowledge that needs to be inferred implicitly. Since the test split is missing, we evaluate models on dev split instead.
  • SocialIQA (SIQA) Sap et al. (2019): Similar to CosQA, this is another MCQ benchmark for commonsense reasoning about social situations that probes emotional and social intelligence in a variety of everyday situations. This dataset has more nuanced commonsense knowledge requirements, which makes it a challenging dataset for our task. Similarly, results on dev split are reported given test is missing.
@@
  • BoolQ Clark et al. (2019): This is a True/False QA dataset consisting of aggregated queries to the Google search engine. Questions in this dataset require complex and difficult entailment-like inference to solve, making it a good set for evaluation. The test split is lacking for this dataset and we can only report results on dev split.
@@
  • Entailer Tafjord et al. (2022): This dataset contains entailment-style statements and corresponding rationales obtained from EntailmentBank dataset Dalvi et al. (2021) and crowdsourcing. The dataset mainly contains science domain statements and tests simple deductive reasoning skills whereby sentences from the premise have to be combined to either support or refute the hypothesis. The test split is also missing for this dataset and we can only evaluate on the dev split.
  • ECQA Aggarwal et al. (2021): This is a human-annotated explanation dataset for CommonsenseQA Talmor et al. (2019). We only use the explanations for the correct choice as the explanations for the incorrect choices are often trivial. It is a complete dataset and, by convention, we use the test split for evaluation.
@@ B.1 Task-Finetuned models
Nie et al. (2020); Liu et al. (2019) This is a strong pre-trained RoBERTa-Large model with corresponding model card on HuggingFace [^3] https://huggingface.co/models called “ynie/roberta-large-snli_mnli_fever_anli_R1_R2_R3-nli”. It is a specifically pre-trained RoBERTa-Large for NLI task and includes the combination of SNLI Bowman et al. (2015), MNLI Williams et al. (2018), FEVER Nie et al. (2019) and ANLI Nie et al. (2020) datasets as the training data. Hence, this incurs a potential data leakage problem, as we also test it on FEVER and ANLI. To some extent, it explains the strong performance of RoBERTa on FEVER and ANLI in Table LABEL:tab:human_results.
The model is used as a classifier in evaluation, and three classes are available. Class 0 corresponds to “entail”, class 1 corresponds to “neutral” and class 2 means “not entail”. In our experiment setting, we only regard class 0 as “Yes” label and combine the remaining two classes to be “No” label.
@@
Tafjord et al. (2022) Entailer is a T5-based model Raffel et al. (2020) trained to answer hypotheses by building proof trees containing chains of reasoning. It can either generate valid premises for a given hypothesis or predict a score for a given premise and hypothesis. We evaluate Entailer-11B (model name “allenai/entailer-11b” on HuggingFace) in our experiments. Similarly, Entailer dataset Dalvi et al. (2021) is in the training set, making this model very competitive when evaluating on the same dataset.
We strictly follow the official implementation of the model to acquire class labels. [^4] https://github.com/allenai/entailment_bank/blob/main/entailer.md The “entailment_verifier” is called to decide if the hypothesis can be implied from the premise. If the answer is “True”, then the class label will be “Yes”. And vice versa.

### v3

@@ 4 Training LLMs for Entailment Verification
In Section 3.2, we observed that open-sourced LLMs lack performance compared to humans and close-sourced models such as GPT-4. From Section 3.1, we know that task-finetuned models perform well on a category when finetuned on data from the same category. Using this insight, we finetune the Flan-T5-xxl model on the train splits of datasets from each category, resulting in using ANLI Nie et al. (2020), RACE Lai et al. (2017), and ECQA Aggarwal et al. (2021) for finetuning. Please refer to Appendix F.1 for more details on our training dataset selection criteria. Next, we describe the finetuning approaches and our key findings.
4.1 Finetuning Formulations
This section describes the two finetuning formulations explored in this work.
Classification
@@ 4.1 Finetuning Formulations
In this approach, for a given premise and hypothesis pair $(p,h)$, we define a weaker hypothesis $h^{\prime}$ as a statement such that the premise $p$ supports $h$ more strongly than $h^{\prime}$. Then, for a given triplet $(p,h,h^{\prime})$, we formulate the ranking task as predicting the hypothesis that is more supported by the premise. Given the triplet $(p,h,h^{\prime})$, we define the margin ranking loss as follows:
@@
where $s(p,h)$ is the entailment score as defined in Equation 1. The key advantage of this formulation over the classification is that ranking, by design, learns a softer decision boundary between the two labels. This can lead to better generalization, especially for contextual QA datasets. Sometimes, the wrong choice can be relatively less favorable compared to the best option in QA instead of being incorrect. As discussed in Section 3.2, this is indicated by the low agreement between human raters. Training to hard-classify the hypothesis for such options can be avoided by ranking them with the best hypothesis (corresponding to the right choice), thus learning a softer classification boundary. We refer to the finetuned model using the ranking objective as “Flan-T5-xxl + Rank”. Please refer to Appendix F.2 for more details on the training data collection process for ranking.
@@ 4.2 Findings
Table LABEL:tab:finetune_results shows the evaluation results on the complete evaluation set (i.e., we use all the data points instead of 100 samples per dataset, which was used in Table LABEL:tab:human_results). For our models, we separately average the results for the datasets already seen in training (namely, ANLI, RACE, and ECQA) and unseen during training into two columns, seen and unseen, respectively. First, we observe that all our finetuned models are consistently better across nine of ten datasets than the baseline Flan-T5-xxl. Finetuning improves 0.07 macro-F1 on average over Flan-T5-xxl. This shows that finetuning is overall beneficial in training the model on the task of entailment verification.
@@
5 Application: Filtering CoT Rationales
Recently, Wang et al. (2023) proposed self-consistency (SC), a decoding technique to improve over chain-of-though (CoT) reasoning Wei et al. (2022) in LLMs, whereby multiple CoT rationales are sampled for a given input instance and a majority voting overall predicted labels is considered as the final prediction. However, generative LLMs can potentially output rationales that are inconsistent Ye and Durrett (2022), i.e., the rationale does not support the corresponding model prediction. Such inconsistency can, in turn, degrade the overall self-consistency results. Evaluating the consistency between an explanation and the corresponding prediction can be framed as an entailment verification (EV) task, as described below. Please refer to Appendix G.2 for examples of consistent and inconsistent CoTs.
Approach
As shown in Figure LABEL:fig:cot_application, we can use a verifier as an intermediate filtering step to filter out the inconsistent rationales before computing the majority vote. For this, we define the generated CoT rationale as the premise and use the QA-to-statement model Chen et al. (2021) as defined in Section 2.1 to convert the question and model’s prediction into a hypothesis. Next, we calculate the entailment score of all the premise-hypothesis pairs using a verifier (Equation 1). Finally, we select the top-$k$ rationales for majority voting, discarding the rest. We set $k=5$ for all our experiments.
@@ 5 Application: Filtering CoT Rationales
In figure LABEL:fig:cot_comparison, we compare the vanilla SC with the filtering+SC approach described above. Following Wang et al. (2023), we compute the average performance of these methods across three MCQ datasets for four different base CoT models (UL2 Tay et al. (2023), Codex-001 Brown et al. (2020), LaMDA-137B Thoppilan et al. (2022), and ChatGPT [^5] ChatGPT refers to the gpt-3.5-turbo-0613 model. OpenAI (2022)). Please refer to Appendix G for details on the datasets and more comparisons with Flan-T5-xxl. We observe that filtering leads to a consistent performance gain over SC across all CoT base models. This demonstrates the advantage of the filtering approach. Next, we find that the improvements are more prominent for weaker base models such as UL2 than the stronger ones (ChatGPT). For instance, filtering UL2 generated-rationales can even achieve comparable performance with vanilla SC over LaMDA-137B. In comparison, the gains for filtering ChatGPT CoTs are $\sim$ 1.7 %. This shows that weaker models are prone to generating inconsistent CoTs and thus benefit more from this approach. But at the same time, even stronger models such as ChatGPT can still benefit from consistency checks. Please refer to Appendix G.2 for examples of filtered CoT rationales.
@@ 6 Related Works
With the advent of general-purpose LLMs Brown et al. (2020); Chung et al. (2022); OpenAI (2023), many prompting strategies have been proposed to generate a natural language reasoning along with the model’s prediction Wei et al. (2022); Zhou et al. (2023); Yao et al. (2023); Huang and Chang (2023). Recently, Ye and Durrett (2022) have found that such generations can sometimes be unreliable due to non-factual and inconsistent reasoning, while Huang et al. (2023) have argued that LLMs struggle to self-correct such issues without external feedback.
Prior works have addressed this limitation by oversampling reasoning chains and marginalizing Wang et al. (2023), using the LLMs itself to recheck their reasoning Madaan et al. (2023); Miao et al. (2023), leveraging external knowledge source to verify factuality Zhao et al. (2023), using deterministic solvers to improve faithfulness Lyu et al. (2023), decomposing the reasoning steps into smaller steps Ling et al. (2023), etc. While the progress is impressive, some of these are either specialized approaches for math-specific datasets or heavily rely on close-sourced LLMs (GPT-3.5, GPT-4, etc.) for verification. In contrast, here we focus on natural language datasets and develop an open-sourced model for the verification task.
Ranking Objective
Prior works have explored the benefits of ranking objective both from a theoretical perspective Narasimhan and Agarwal (2013) and specifically for NLP classification tasks Li et al. (2019); Briakou and Carpuat (2020). Our ranking-based objective is inspired by these works that find that margin-based loss can be beneficial in training plausibility estimation models.
@@ 7 Conclusion
We studied the EV problem in the context of LLMs. Specifically, we sourced datasets across three different categories (NLI, contextual QA, and rationales) and analyzed human and LLM performance on the benchmark. We found that LLMs are better than humans in complex reasoning, while humans are more consistent on simpler reasoning tasks. We also explored different training objectives to finetune open-sourced LLMs for EV. Our finetuned model outperforms GPT-3.5, baseline Flan-T5-xxl, and is comparable to GPT-4. Finally, we applied the EV model to filter out inconsistent model-generated CoTs in self-consistency decoding, achieving improvements over the baseline self-consistency approach.
@@
Even though our work demonstrates exciting results on entailment verification tasks by finetuning LLMs, several limitations can be potentially improved.
Data Processing
Our strategy to convert a QA pair into a statement using the QA-to-statement converter model can have errors that can cascade both in the evaluation dataset and our fine-tuned models. Better strategies for this pipeline would help with data quality.
Finetuning
We only tried encoder-decoder-based models for finetuning. However, other models with different architectures (like decoder-only) can also be considered in the future. For computing the entailment score in Equation 1, we only considered the probability of “Yes” and “No” tokens in the entire vocabulary. Other alternative expressions like “YES”/“NO”, “True”/“False”, etc., can also be considered to make the score more robust. Finally, our training objective outputs entailment scores instead of directly generating answers. Answer generation as a training objective can be more robust since it is a stricter objective than our scoring technique.
Potential Risks
We using existing datasets, with some post-processing, to train our models. Thus, any issues in the existing dataset in terms of bias, toxicity, etc., can potentially affect our model. Also, since we use a pretrained checkpoint, we also inherit any existing biases in the baseline model. The model is trained to always output scores, even if the data is well outside the training distribution. This is an existing issue with most NLP models and can be mitigated by additional checks for domain shift.
@@ Acknowledgments
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200006, the Defense Advanced Research Projects Agency with award HR00112220046, and NSF IIS 2048211. The views and conclusions contained herein are those of the authors. They should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. We would like to thank all the USC INK research lab collaborators for their constructive feedback on the work.
@@ Appendix A Evaluation Datasets
NLI is an obvious choice of data source as it is a more general case of the entailment verification problem. While converting an NLI dataset for our task, we merge the neutral and contradict labels to the not support label. [^6] We note that merging the labels for the NLI datasets in this benchmark is valid, since the underlying meaning of the ‘neutral’ label is ‘no sufficient evidence to support’ in all three datasets. NLI datasets that don’t follow this label definitions cannot be modified in this manner. We use the following NLI datasets in our benchmark:
  • WaNLI Liu et al. (2022): This is a new NLI dataset built using worker and AI collaboration. This challenging dataset improves over the existing NLI dataset MultiNLI Williams et al. (2018). The test split is used when doing the evaluation.
  • FEVER Nie et al. (2019): This is a modification of the original FEVER dataset Thorne et al. (2018) in which the claim is paired with textual evidence from Wikipedia to convert it into an NLI format dataset. This pairing uses existing state-of-the-art evidence extraction systems to find relevant evidence for each claim. Premises in this dataset typically contain multiple sentences, which is one of our focus areas. As the test split is unavailable, we report results on the dev split for evaluation.
@@
Next, we consider QA datasets where the task is to answer a question based on a given context and some options. We use an off-the-shelf QA-to-statement converter model Chen et al. (2021) to generate a hypothesis statement for each question option pair. Then, the hypothesis corresponding to the correct choice is marked as “support”, while the rest are marked as “not support” to create the entailment verification dataset. We depict this process in Figure LABEL:fig:contextual_qa. The green box is the valid hypothesis and the red ones are the invalid hypothesis corresponding to the given context. Overall, we include the following datasets from this category:
  • Cosmos QA (CosQA) Huang et al. (2019): This dataset contains multiple-choice questions (MCQs) that require an understanding of commonsense-based reading comprehension to answer a question. The key challenge in this dataset is understanding people’s everyday narratives described in the context that can have some missing commonsense knowledge that needs to be inferred implicitly. Since the test split is missing, we evaluate models on the dev split instead.
  • SocialIQA (SIQA) Sap et al. (2019): Similar to CosQA, this is another MCQ benchmark for commonsense reasoning about social situations that probes emotional and social intelligence in a variety of everyday situations. This dataset has more nuanced commonsense knowledge requirements, which makes it a challenging dataset for our task. Similarly, results on the dev split are reported, given that the test is missing.
@@
  • BoolQ Clark et al. (2019): This is a True/False QA dataset consisting of aggregated queries to the Google search engine. Questions in this dataset require complex and difficult entailment-like inference to solve, making it a good set for evaluation. The test split is lacking for this dataset, and we can only report results on the dev split.
@@
  • Entailer Tafjord et al. (2022): This dataset contains entailment-style statements and corresponding rationales obtained from EntailmentBank dataset Dalvi et al. (2021) and crowdsourcing. The dataset mainly contains science domain statements and tests simple deductive reasoning skills whereby sentences from the premise must be combined to either support or refute the hypothesis. The test split is also missing for this dataset, and we can only evaluate it on the dev split.
  • ECQA Aggarwal et al. (2021): This is a human-annotated explanation dataset for CommonsenseQA Talmor et al. (2019). We only use the explanations for the correct choice as the explanations for the incorrect choices are often trivial. It is a complete dataset and by convention, we use the test split for evaluation.
@@ B.1 Task-Finetuned models
Nie et al. (2020); Liu et al. (2019) This is a strong pre-trained RoBERTa-Large model with the corresponding model card on HuggingFace [^7] https://huggingface.co/models Wolf et al. (2020b) called “ynie/roberta-large-snli_mnli_fever_anli_R1_R2_R3-nli”. It is a specifically pre-trained RoBERTa-Large for NLI task and includes the combination of SNLI Bowman et al. (2015), MNLI Williams et al. (2018), FEVER Nie et al. (2019) and ANLI Nie et al. (2020) datasets as the training data. Hence, this incurs a potential data leakage problem, as we also test it on FEVER and ANLI. To some extent, it explains the strong performance of RoBERTa on NLI datasets in Table LABEL:tab:human_results.
The model is used as a classifier in evaluation, and three classes are available. Class 0 corresponds to “entail”, class 1 corresponds to “neutral” and class 2 means “not entail”. In our experiment setting, we only regard class 0 as the “Yes” label and combine the remaining two classes to be the “No” label.
@@
Tafjord et al. (2022) Entailer is a T5-based model Raffel et al. (2020) trained to answer hypotheses by building proof trees containing chains of reasoning. It can either generate valid premises for a given hypothesis or predict a score for a given premise and hypothesis. We evaluate Entailer-11B (model name “allenai/entailer-11b” on HuggingFace) in our experiments. Similarly, Entailer dataset Dalvi et al. (2021) is in the training set, making this model very competitive when evaluating the same dataset.
We strictly follow the official implementation of the model to acquire class labels. [^8] https://github.com/allenai/entailment_bank/blob/main/entailer.md The “entailment_verifier” is called to decide if the hypothesis can be implied from the premise. If the answer is “True”, then the class label will be “Yes” and vice versa.